1. A small land rover, designed to explore the Martian surface, weighs 1.8 x 103 N on Earth. How much will it weigh on Mars, where the gravitational field strength there is 3.3 N/kg?
  2. A single cable is used to hang a 134 kg sign above a store entrance. What is the tension force acting on the cable?
  3. If a net force of 68.0 N acts on a 4.05 kg mass, what is its acceleration?
  4. A model rocket has a mass of 0.435 kg. It is mounted vertically and launches directly up with an thrust force of 13.1 N.

A) What is the weight of the rocket?

B) Determine the net force on the rocket as it is launched.



1
Expert's answer
2021-04-19T17:11:39-0400

1. We can find the mass of the rover in kilograms from its weight on Earth:


"W_E=mg_E,\\\space\\
m=\frac{W_E}{g_E},\\
W_M=mg_M=W_E\frac{g_M}{g_E}=5345\text{ N}."

2. The tension can be found from the weight of the body:


"T=W=mg=1313\text{ N}."

3. According to Newton's second law:


"A=\frac Fm=16.8\text{ m/s}^2."

4. A) The weight is:


"W=mg=4.263\text{ N}."

B) The net force:


"F_\text{net}=F-W=13.1-4.263=8.837\text{ N}."